OpenStudy (neonclawzz):

I have part A done but i need help with part B. Will medal! Given he function f(x) = 4(2)^x, Section A is from x = 1 to x = 2 and Section B is from x = 3 to x = 4. Part A: Find the average rate of change of each section. Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other

OpenStudy (neonclawzz):

My answer- Part a) section A f(average)= 4(2)^2 - 4(2)^1 ____________ = 8 2-1 section B f(average)= 4(2)^4 - 4(2)^3 ____________ = 32 4-3 Part b) The rate of change is 4 times greater between section B than between section A because.....(that's all I have)

OpenStudy (loser66):

32= ?*8

OpenStudy (neonclawzz):

whoops sorry i left some out of the question. this is whats missing- Explain why one rate of change is greater than the other

OpenStudy (neonclawzz):

I have to explain but i'm not sure how

OpenStudy (loser66):

and you get the right answer, just 4 times. Dat sit

OpenStudy (loser66):

And you need the explanation part, right?

OpenStudy (loser66):



OpenStudy (loser66):

the exponential function is monotonously increasing. That explains why the rate of section B is greater than it is on section A

OpenStudy (neonclawzz):

Okay thank you so much :>